# The Born approximation for the fixed energy Calderón problem

### v1

@@ 1.3. Numerical reconstruction
In this subsection, we show the capabilities of the Born approximation as an effective tool to approximate the potential; numerical methods based on this strategy are described in [BCMM24] in the case $\kappa=0$. The key remark is that the Fréchet differential $\mathrm{d}\Phi^{\kappa}_{0}$ coincides with the differential of $\Phi^{0}$ around the constant function $-\kappa\in\mathbb{R}\setminus\operatorname{Spec}_{H^{1}_{0}}{\Delta}$, and that $\Lambda_{q,0}=\Lambda_{q+\kappa,\kappa}$. Therefore, for such $\kappa$, given $q\in L^{\infty}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ such that 00 is not a Dirichlet eigenvalue of $-\Delta+q$, one should have:
@@
When $\kappa$ is chosen appropriately, so that $q+\kappa$ is small in some norm, one can expect this function to be well-approximated by $(q+\kappa)^{\mathrm{B}}_{\kappa}$. Recall that, by 2, the Born approximation $(q+\kappa)^{\mathrm{B}}_{\kappa}$ is constructed solely in terms of $\operatorname{Spec}(\Lambda_{q,0})$ (the dependence on $\kappa$ being explicit). A possible choice of $\kappa$ that can be implemented numerically consists in ensuring that $\int_{\mathbb{B}^{d}}(q+\kappa)^{\mathrm{B}}_{\kappa}\mathrm{d}x=\mathcal{F}(q+\kappa)^{\mathrm{B}}_{\kappa}(0)$, which is a quantity that depends only on $\operatorname{Spec}(\Lambda_{q,0})$ and $\kappa$, is the smallest possible. It is therefore natural to choose $\kappa=\kappa_{*}$ defined by the equation
@@
In the caption of each plot, we include the $L^{1}(\mathbb{B}^{3})$ norm of the difference of the functions plotted. Notice that in all eight plots, except for Fig. 2 (A) where the Gibbs phenomenon is too strong, $(q+\kappa)-(q+\kappa)^{\mathrm{B}}_{\kappa}$ is a continuous function outside $x=0$ that vanishes at $\mathbb{S}^{2}$, as stated in 3.
@@
Figure 1. Plots of $q(x)=2-\chi_{(\frac{1}{3},\frac{2}{3})}(\left\lvert x\right\rvert)$ (blue) and $(q+\kappa)^{\mathrm{B}}_{\kappa}-\kappa$ (orange).
@@
Figure 3. Plots of $q(x)=2-\chi_{(\frac{1}{3},\frac{2}{3})}(\left\lvert x\right\rvert)$ (blue) and $q^{\mathrm{B}}_{\kappa}$ (orange).
Figure 4. Plots of $q_{i}$ (solid) of $q_{3}(x)\coloneqq 3$, $q_{2}(x)\coloneqq q_{3}(x)-\chi_{(0,\frac{2}{3})}(\left\lvert x\right\rvert)$, $q_{1}(x)\coloneqq q_{2}(x)-\chi_{(0,\frac{1}{3})}(\left\lvert x\right\rvert)$ and their respective Born approximations $q_{i,\kappa}^{\mathrm{B}}$ (dashed) at $\kappa=-1$.
The Julia code used to produce these plots can be found in [MMSM24].
@@ 2. Radial Dirichlet-to-Neumann maps
This problem has a unique solution provided that $q\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$, with $p$ admissible, and $\kappa$ is not a Dirichlet eigenvalue of $-\Delta+q$. In addition, the radiality of $q$ implies that
@@
is an invariant subspace of $-\Delta+q$.  [^2] One uses the standard identification $(v\otimes f)(x)=v_{\mathrm{rad}}(|x|)f(\widehat{x})$, where $\widehat{x}:=x/|x|$ for every $x\in\mathbb{R}^{d}\setminus\{0\}$. More precisely, the following result holds.
@@
Let $p$ be admissible and $q\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d})$. Also, let $\ell\in\mathbb{N}_{0}$,
@@
(2.5) $-\frac{1}{r^{d-1}}\frac{\mathrm{d}missing}{\mathrm{d}r}\left(r^{d-1}\frac{\mathrm{d}missing}{\mathrm{d}r}b_{\ell}(r,\kappa)\right)+\left(\frac{\ell(\ell+d-2)}{r^{2}}+q_{0}(r)-\kappa\right)b_{\ell}(r,\kappa)=0,\quad r\in(0,1),$
with $b_{\ell}(1,\kappa)=1$. The change of variables $w_{\ell}(t):=b_{\ell}(e^{-t},\kappa)e^{-\nu_{d}t}$ transforms (2.5) into a one-dimensional Schrödinger equation on the half-line:
@@
    In other words, the spectrum of $\Lambda_{q,\kappa}$ consists of the eigenvectors:
@@
    When $\kappa\in\mathbb{R}$, it is possible to give a meaningful definition of the eigenvalues $\lambda_{\ell}[q,\kappa]$ for every $\ell$ large enough, even when $\kappa$ is a Dirichlet eigenvalue of $-\Delta+q$, see [DMMN24, Definition 2.4]. In [DMMN24, Lemma 2.1] it is shown that (2.1) has a unique solution of the form (2.2) that belongs to $H^{1}_{\rm loc}(\mathbb{B}^{d}\setminus\{0\})\cap L^{2}_{\nu_{d}-1/2}(\mathbb{B}^{d})$.
@@
Let $p$ be admissible and $\kappa\in\mathbb{C}\setminus\operatorname{Spec}_{H^{1}_{0}}(-\Delta)$. If $q\in L^{p}(\mathbb{B}^{d},\mathbb{R})$ is such that $\kappa$ is not a Dirichlet eigenvalue of $-\Delta+q$ then, for every $f,g\in L^{2}(\mathbb{S}^{d-1})$,
@@
(2.9) $\left\{\begin{array}[]{rll}(-\Delta+q-\kappa)v&=0,&\text{ in }\mathbb{B}^{d},\\ v&=f,&\text{ on }\mathbb{S}^{d-1},\end{array}\right.\qquad\left\{\begin{array}[]{rll}-\Delta u-\overline{\kappa}u&=0,&\text{ in }\mathbb{B}^{d},\\ u&=g,&\text{ on }\mathbb{S}^{d-1}.\end{array}\right.$
@@
where $u$ is the same as in (2.9), and $w$ satisfies that $(-\Delta-\kappa)w=0$ on $\mathbb{B}^{d}$ and $w|_{\mathbb{S}^{d-1}}=g$. Note that the previous expression ensures that $\mathrm{d}\Phi^{\kappa}_{0}(q)$ is a bounded linear operator on $L^{2}(\mathbb{S}^{d-1})$. Taking $f=g=Y_{\ell}$ with $Y_{\ell}\in\mathfrak{H}_{\ell}^{d}$ normalized, by Lemma 2.1 and (2.6) we find that $w(x)=b_{\ell}(|x|,\kappa)Y_{\ell}(\hat{x})$ and $\overline{u(x)}=b_{\ell}(|x|,\kappa)\overline{Y_{\ell}(\hat{x})}$, where $b_{\ell}$ satisfies (2.3), since we are in the case $q=0$. Therefore, we have
@@
where $v$ satisfies (2.9) with $f=Y_{\ell}$. Let $h(x):=v(x)-b_{\ell}(|x|,\kappa)Y_{\ell}(\widehat{x})$. Inserting this in the previous identity and using that $Y_{\ell}$ is normalized, yields
@@
Denoting by $\mathcal{R}_{q}(\kappa)$ the Dirichlet resolvent $(-\Delta+q-\kappa)^{-1}$, we can write $h=-\mathcal{R}_{q}(\kappa)(qg_{\kappa})$ with $g_{\kappa}(x):=b_{\ell}(|x|,\kappa)Y_{\ell}(\widehat{x})$. Since $-\Delta+q$ and $\mathcal{R}_{q}(\kappa)$ leave invariant the subspace $\mathcal{E}_{\ell}$ we have
@@ 3. Fourier representation formulas
In this section, we prove 1. Recall that $\mathcal{P}_{\ell,d}$ denotes the orthogonal projection of $L^{2}(\mathbb{S}^{d-1})$ onto $\mathfrak{H}_{\ell}^{d}$.
@@
$=\sum_{\ell=0}^{\infty}\int_{0}^{R}f_{0}(r)r^{d-1}\left\langle\overline{\mathcal{P}_{\ell,d}e_{r\zeta_{1}}},\mathcal{P}_{\ell,d}e_{r\zeta_{2}}\right\rangle_{L^{2}(\mathbb{S}^{d-1})}\mathrm{d}r.$
The exchange of series and integral is justified by Fubini’s Theorem and the bound
@@
The definition of the Legendre polynomial $P_{\ell,d}$ (see Appendix B) implies that the complex phases of all its terms align when we evaluate it at $i\mathbb{R}$, therefore
@@
Putting all these bounds together we see that we can indeed apply the dominated convergence theorem in order to exchange limit and series. The proof is then finished by the convergence of $f*\eta_{\varepsilon}$ to $f$ in $\mathcal{E}^{\prime}(\mathbb{R}^{d})$.
@@ 4.1. Weyl-Titchmarsh functions and DtN maps
(4.1) $\left\{\begin{array}[]{rll}(-\partial_{t}^{2}+Q(t))v_{z}(t)&=-z^{2}v_{z}(t),&\text{ in }\mathbb{R}_{+},\\ v_{z}(0)&=1,&\end{array}\right.$
@@
(4.2) $Q\in A^{1}(\mathbb{R}_{+})\quad\iff\quad{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|Q\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}:=\sup_{x\in\mathbb{R}_{+}}\int_{x}^{x+1}|Q(t)|\,\mathrm{d}t<\infty,$
@@
When $02p-d>$ 0 one can improve this to:
@@
which corresponds via (4.6) to the constant potential $-\kappa$. When $\kappa=0$, it is not difficult to check that its Weyl-Titchmarsh function equals $m_{0}(-z^{2})=-z$. When $\kappa\neq 0$, $m_{Q_{\kappa}}$ can be explicitly computed as well.
@@
Suppose $\kappa\in\mathbb{C}\setminus\{0\}$. For every $z\geq 0$ such that $J_{z}(\sqrt{\kappa})\neq 0$, [^3] This condition is always fulfilled when $\kappa\in\mathbb{C}\setminus\mathbb{R}_{+}$, since all the roots of $J_{z}$ are real as soon as $z>-1$. the problem
@@
The condition $u_{z}(0,\kappa)=1$ translates into $w_{z}(1,\kappa)=1$; and this forces
@@
Identity (4.10) follows by taking into account that $\operatorname{Spec}_{H^{1}_{0}}(-\Delta|_{\mathcal{E}_{\ell}})$, the spectrum of the restriction of the Dirichlet Laplacian $-\Delta$ on $\mathbb{B}^{d}$ to functions with angular component in $\mathfrak{H}^{d}_{\ell}$, is precisely the set of solutions of $\varphi_{\ell}(\sqrt{\kappa})=0$. Finally, the calculation of the spectrum of $\Lambda_{0,\kappa}$ is obtained by (4.9) and the last identity of Lemma 4.1. ∎
@@ 4.2. Fixed energy $A_{\kappa}$-amplitude
(4.12) $z_{Q}:=2e({\left|\kern-1.07639pt\left|\kern-1.07639pt\left|Q\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}+\sqrt{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|Q\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}}),$
@@
This follows from the fact that $J_{z}(\sqrt{\kappa})$ has no non-negative zeros for $\kappa$ fixed if $\kappa<0$. If $\kappa>0$, then the largest zero $z_{L}$ satisfies $j_{z_{L},1}=\sqrt{\kappa}$, and there are no zeros when $z$ is in the region determined by the condition $j_{z,1}>\sqrt{\kappa}$. Since from [LM83, Equation (2.4)] we have the lower bound $j_{z,1}\geq z+j_{0,1}$ when $z\geq 0$, it follows that $J_{z}(\sqrt{\kappa})$ has no zeros if $z+j_{0,1}>\sqrt{\kappa}$ and $z\geq 0$. In other words, non-negative zeros are contained in the interval $[0,\sqrt{\kappa}-j_{0,1}]$ when $\sqrt{\kappa}\geq j_{0,1}$, and there are no zeros in $[0,\infty)$ if $\sqrt{\kappa}<j_{0,1}$. By [RS00, Theorem 1.3] we conclude the existence of $A_{Q_{\kappa}}\in L^{1}_{z_{\kappa}}(\mathbb{R}_{+})$ where $z_{\kappa}$ is defined in the statement. ∎
@@
    (4.15) $|A_{Q,\kappa}(t)+Q_{\kappa}(t)-Q(t)|\leq C_{\kappa}e^{t(2\alpha_{Q}(t)+|\kappa|)}\int_{0}^{t}|Q(s)-Q_{\kappa}(s)|\,\mathrm{d}s.$
@@
    If $Q\in\mathcal{C}^{m}(\mathbb{R}_{+})$ for some $m\in\mathbb{N}_{0}$, then $A_{Q,\kappa}-Q$ belongs to $\mathcal{C}^{m+1}(\mathbb{R}_{+})$.

### v2

@@ 1.3. Numerical reconstruction
In this subsection, we show the capabilities of the Born approximation as an effective tool to approximate the potential; numerical methods based on this strategy are described in [BCMM24] in the case $\kappa=0$. The key remark is that the Fréchet differential $\mathrm{d}\Phi^{\kappa}_{0}$ coincides with the differential of $\Phi^{0}$ around the constant function $\kappa\in\mathbb{R}\setminus\operatorname{Spec}_{H^{1}_{0}}{\Delta}$, and that $\Lambda_{q,0}=\Lambda_{q+\kappa,\kappa}$. Therefore, for such $\kappa$, given $q\in L^{\infty}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ such that 0 is not a Dirichlet eigenvalue of $\Delta+q$, one should have:
@@
When $\kappa$ is chosen appropriately, so that $q+\kappa$ is small in some norm, one can expect this function to be well-approximated by $(q+\kappa)^{\mathrm{B}}_{\kappa}$. Recall that, by 1, the Born approximation $(q+\kappa)^{\mathrm{B}}_{\kappa}$ is constructed solely in terms of $\operatorname{Spec}(\Lambda_{q,0})$ (the dependence on $\kappa$ being explicit). A possible choice of $\kappa$ that can be implemented numerically consists in ensuring that $\int_{\mathbb{B}^{d}}(q+\kappa)^{\mathrm{B}}_{\kappa}\mathrm{d}x=\mathcal{F}(q+\kappa)^{\mathrm{B}}_{\kappa}(0)$, which is a quantity that depends only on $\operatorname{Spec}(\Lambda_{q,0})$ and $\kappa$, is the smallest possible. It is therefore natural to choose $\kappa=\kappa_{*}$ defined by the equation
@@
In the caption of each plot, we include the $L^{1}(\mathbb{B}^{3})$ norm of the difference of the functions plotted. Notice that in all eight plots, except for Fig. 2 (A) where the Gibbs phenomenon is too strong, $(q+\kappa)-(q+\kappa)^{\mathrm{B}}_{\kappa}$ is a continuous function outside $x=0$ that vanishes at $\mathbb{S}^{2}$, as stated in 2.
The Julia code used to produce these figures can be found in [MMSM24].
@@
[…]
@@
[…]
@@ 2. Radial Dirichlet-to-Neumann maps
This problem has a unique solution provided that $q\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$, with $p$ admissible, and $\kappa$ is not a Dirichlet eigenvalue of $\Delta+q$. In addition, the radiality of $q$ implies that
@@
[…]
@@
Let $p$ be admissible and $q\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d})$. Let $\ell\in\mathbb{N}_{0}$
@@
(2.5) $\frac{1}{r^{d-1}}\frac{\mathrm{d}}{\mathrm{d}r}\left(r^{d-1}\frac{\mathrm{d}}{\mathrm{d}r}b_{\ell}(r,\kappa)\right)+\left(\frac{\ell(\ell+d-2)}{r^{2}}+q_{0}(r)-\kappa\right)b_{\ell}(r,\kappa)=0,\quad r\in(0,1),$
with $b_{\ell}(1,\kappa)=1$. The Liouville change of variables $w_{\ell}(t):=b_{\ell}(e^{-t},\kappa)e^{-\nu_{d}t}$ (see, for instance, [BR89, Chapter 10.9]) transforms (2.5) into a one-dimensional Schrödinger equation on the half-line:
@@
    In other words, the spectrum of $\Lambda_{q,\kappa}$ consists of the eigenvalues:
@@
    When $\kappa\in\mathbb{R}$, it is possible to give a meaningful definition of the eigenvalues $\lambda_{\ell}[q,\kappa]$ for every $\ell$ large enough, even when $\kappa$ is a Dirichlet eigenvalue of $\Delta+q$, see [DMMN24, Definition 2.4]. In [DMMN24, Lemma 2.1] it is shown that (2.1) has a unique solution of the form (2.2) that belongs to $H^{1}_{\rm loc}(\mathbb{B}^{d}\setminus\{0\})\cap L^{2}_{\nu_{d}-1/2}(\mathbb{B}^{d})$.
@@
Let $p$ be admissible and $\kappa\in\mathbb{C}\setminus\operatorname{Spec}_{H^{1}_{0}}(-\Delta)$. If $q\in L^{p}(\mathbb{B}^{d},\mathbb{R})$ is such that $\kappa$ is not a Dirichlet eigenvalue of $\Delta+q$ then, for every $f,g\in L^{2}(\mathbb{S}^{d-1})$,
@@
[…]
Proof.
First take $f,g\in H^{1/2}(\mathbb{S}^{d-1})$. Green’s formula, when applied to a solution $v$ of (2.9) and any function $u\in H^{1}(\mathbb{B}^{d})$ with $u|_{\mathbb{S}^{d-1}}=g$, gives:
$\left\langle g,\Lambda_{q,\kappa}f\right\rangle_{L^{2}(\mathbb{S}^{d-1})}=\int_{\mathbb{B}^{d}}(q(x)-\kappa)\,\overline{u(x)}\,v(x)\,\mathrm{d}x+\int_{\mathbb{B}^{d}}\overline{\nabla u(x)}\,\nabla v(x)\,\mathrm{d}x.$
This yields, as soon as $q$ is real-valued,
$\left\langle g,\Lambda_{q,\kappa}f\right\rangle_{L^{2}(\mathbb{S}^{d-1})}=\left\langle\Lambda_{q,\overline{\kappa}}g,f\right\rangle_{L^{2}(\mathbb{S}^{d-1})}.$
Formula (2.8) then follows by particularizing those identities to $q=0$ and subtracting. In addition, (2.8) shows that $\Lambda_{q,\kappa}-\Lambda_{0,\kappa}$ is bounded on $L^{2}(\mathbb{S}^{d-1})$: for some $C=C_{q,d}>0$,
$\left\langle g,(\Lambda_{q,\kappa}-\Lambda_{0,\kappa})f\right\rangle_{L^{2}(\mathbb{S}^{d-1})}\leq\left\lVert q\right\rVert_{L^{\infty}(\mathbb{B}^{d})}\left\lVert u\right\rVert_{H^{1/2}(\mathbb{B}^{d})}\left\lVert v\right\rVert_{H^{1/2}(\mathbb{B}^{d})}\leq C\left\lVert f\right\rVert_{L^{2}(\mathbb{S}^{d-1}}\left\lVert g\right\rVert_{L^{2}(\mathbb{S}^{d-1})}$
and therefore (2.8) holds for $f,g\in L^{2}(\mathbb{S}^{d-1})$ as claimed. ∎
@@
[…]
@@
[…]
@@
[…]
@@ 3. Fourier representation formulas
In this section we prove an explicit formula to reconstruct a distribution from its moments. The result below can be viewed as an explicit formula for $(\mathrm{d}\Phi^{\kappa}_{0})^{-1}|_{\mathcal{E}_{\mathrm{rad}}^{\prime}(\mathbb{R}^{d})}$. In particular, it implies that, as soon as $q^{\mathrm{B}}_{\kappa}\in\mathcal{B}_{d}$ exists, it is uniquely determined by (1.12) in $\mathcal{B}_{d}$.
Theorem 3.1.
where the series are absolutely convergent.
Note that taking the limit $\kappa\to 0$ in (3.1) recovers the representation for $\kappa=0$ that was obtained in [BCMM22, DMMN24], see (1.23).
Proof.
@@
$=\sum_{\ell=0}^{\infty}\int_{0}^{R}f_{0}(r)r^{d-1}\left\langle\overline{\mathcal{P}_{\ell,d}e_{r\zeta_{1}}},\mathcal{P}_{\ell,d}e_{r\zeta_{2}}\right\rangle_{L^{2}(\mathbb{S}^{d-1})}\mathrm{d}r,$
where $\mathcal{P}_{\ell,d}$ denotes the orthogonal projection of $L^{2}(\mathbb{S}^{d-1})$ onto $\mathfrak{H}_{\ell}^{d}$. The exchange of series and integral is justified by Fubini’s Theorem and the bound:
@@
The definition of the Legendre polynomial $P_{\ell,d}$ (see Appendix B) implies that the complex phases of all its terms align when we evaluate it at $i\mathbb{R}$; therefore
@@
Putting all these bounds together, we see that we can indeed apply the dominated convergence theorem in order to exchange limit and series. The proof is then finished by the convergence of $f*\eta_{\varepsilon}$ to $f$ in $\mathcal{E}^{\prime}(\mathbb{R}^{d})$. ∎
@@ 4.1. Weyl-Titchmarsh functions and DtN maps
[…]
@@
[…]
@@
When 2pd>02p-d>0 one can improve this to:
@@
which corresponds via (4.6) to the constant potential $\kappa$. When $\kappa=0$, it is not difficult to check that its Weyl-Titchmarsh function equals $m_{0}(-z^{2})=-z$. When $\kappa\neq 0$, $m_{Q_{\kappa}}$ can be explicitly computed as well.
@@
Suppose $\kappa\in\mathbb{C}\setminus\{0\}$. For every $z\geq 0$ such that $J_{z}(\sqrt{\kappa})\neq 0$, [^3] This condition is always fulfilled when $\kappa\in\mathbb{C}\setminus\mathbb{R}_{+}$, since all the roots of $J_{z}$ are real as soon as z>1z>-1. the problem
@@
The condition $u_{z}(0,\kappa)=1$ translates into $w_{z}(\sqrt{\kappa},\kappa)=1$; and this forces
@@
Identity (4.10) follows by taking into account that $\operatorname{Spec}_{H^{1}_{0}}(-\Delta|_{\mathcal{E}_{\ell}})$, the spectrum of the restriction of the Dirichlet Laplacian $\Delta$ on $\mathbb{B}^{d}$ to functions with angular component in $\mathfrak{H}^{d}_{\ell}$, is precisely the set of solutions of $\varphi_{\ell}(\sqrt{\kappa})=0$. Finally, the calculation of the spectrum of $\Lambda_{0,\kappa}$ is obtained by (4.9) and the last identity of Lemma 4.1. ∎
@@ 4.2. Fixed energy $A_{\kappa}$-amplitude
[…]
@@
This follows from the fact that $J_{z}(\sqrt{\kappa})$ has no non-negative zeros for $\kappa$ fixed if $\kappa<0$. If $\kappa>0$, then the largest zero $z_{L}$ satisfies $j_{z_{L},1}=\sqrt{\kappa}$, and there are no zeros when $z$ is in the region determined by the condition $j_{z,1}>\sqrt{\kappa}$. Since from [LM83, Equation (2.4)] we have the lower bound $j_{z,1}\geq z+j_{0,1}$ when $z\geq 0$, it follows that $J_{z}(\sqrt{\kappa})$ has no zeros if $z+j_{0,1}>\sqrt{\kappa}$ and $z\geq 0$. In other words, non-negative zeros are contained in the interval $0,\sqrt{\kappa}-j_{0,1}]$ when $\sqrt{\kappa}\geq j_{0,1}$, and there are no zeros in $0,\infty)$ if $\sqrt{\kappa}<j_{0,1}$. By [RS00, Theorem 1.3] we conclude the existence of $A_{Q_{\kappa}}\in L^{1}_{z_{\kappa}}(\mathbb{R}_{+})$ where $z_{\kappa}$ is defined in the statement. ∎
@@
    (4.15) $|A_{Q,\kappa}(t)+Q_{\kappa}(t)-Q(t)|\leq C_{\kappa}(1+\alpha_{Q}(t))e^{t(\alpha_{Q}(t)+|\kappa|)}\int_{0}^{t}|Q(s)-Q_{\kappa}(s)|\,\mathrm{d}s.$
@@
    If $Q\in\mathcal{C}^{m}([0,\infty))$ for some $m\in\mathbb{N}_{0}$, then $A_{Q,\kappa}-Q$ belongs to $\mathcal{C}^{m+1}([0,\infty))$.